OpenStudy (anonymous):

a paper weight is dropped from the roof of a block of multi storey flats. Each storey being 3 metre high. it passes the ceiling of the 20th floor at 30m/s. if g=10 , how many storeys does the building has

OpenStudy (anonymous):

Hint v = v0 + gt v0 = 0

OpenStudy (anonymous):

v - velocity g - gravitational constant t - time v0- initial velocity

OpenStudy (anonymous):

plzz explain

OpenStudy (anonymous):

see my conclusions Let the number of floors in the building be x heght of building = 3x distance travelled = 3x20=60 initial velocity (U)=0 final velocity (v)=30 m/s^2 height building after 20th floor = x-60

OpenStudy (anonymous):

No you have to find the distance travelled by paper not anything else. When you will get the time of flight using the first formula all the way down from top to the 20th story. Next thing to do would be to find how many meters it passed down fro the top. General formula of kinematics for motion in one dimension, with constant acceleration z = z0 + v0t + gt^2/2 z0 - initial distance traveled by it from the top, it is 0 because it is at the top at the beginning v0 - as I mentioned above is also zero so we get eventually z = gt^2/2 which is the asked distance, dividing it by 3, you will get the number of stories it passed, And the answer would Number of stories of the building = z/3+20

OpenStudy (anonymous):

When anything concerns the motion in one dimension, it is enough to know only to formulas of kinematics z = z0 + v0t + at^2/2 and v = v0+at Of course, acceleration must be constant, otherwise the equation becomes more complicated.

OpenStudy (anonymous):

Instead of z there could , x or y, depending on direction of motion.

OpenStudy (anonymous):

Do you have any questions?

OpenStudy (anonymous):

Your conclusion is wrong in one place distance travelled = 3x20=60 We don't know the number of stories it passed down the top, it is the data we asked for. So it is incorrect, I don't get why you decided to use it.

OpenStudy (anonymous):

\[v ^{2}=u ^{2}+2gh\] since the weight is dropped u=0. So \[v ^{2}=2gh\] or \[h=v ^{2}\div2g\] here v=30 m/s and \[g=10 m/s ^{2}\] so, h=45m i.e 15 floors. therefore the building is 20+15=35 floors high.